# Floods do not sink prices, historical memory does: How flood risk impacts the Italian housing market

### v2

@@ 3 Econometric design
To quantify the influence of flood risk on the Italian housing market, we have addressed three separate research questions: (1) Do we observe changes in home prices after specific flood events?; (2) Does flood risk affect home prices? Is this influenced by repeated exposure to floods (historical memory)?; (3) How do results change across socio-economic groups? In the following, we describe how we address all these questions by building on a baseline hedonic regression.
@@ 3.1 Baseline hedonic regression
In our formulation, $\alpha_{k}$ captures all time-invariant home price differences due to being located in spatial unit $k$. We consider different definitions of spatial units, at different spatial aggregation levels. The widest definition of $k$ is the municipality, with a median area of 22 km2. Especially for large municipalities, it is unlikely that this fixed effect captures enough difference in location, so we consider two further narrower definitions of $\alpha_{k}$. These are the OMI microzone and the census tract, respectively featuring a median area of 1.5 km2 and 0.03km2.252525The OMI microzone is the standard spatial unit of the Osservatorio del Mercato Immobiliare (OMI), the housing market observatory of the Italian Tax Office. It roughly corresponds to a neighborhood, although in rural areas it may contain larger portions of territory. Instead, census tracts correspond to blocks in cities and to villages (“frazioni”) in rural areas. We use the latest available census tracts, namely from the 2011 census. The inclusion of the $\alpha_{k}$ term completely captures the spatial correlation existing in our data, see Supplementary Section B.2. This confirms that our hedonic regression approach is adequate and spatial econometric approaches are not necessary.
@@
3.2 Effect of flood events
To estimate the impact that specific flood events had on home price (research question 1) we built on Equation (2) to create a hedonic difference-in-difference specification. Following Ortega and Taṣpınar (2018), we disaggregated the $R_{i}$ indicator into two categories, homes in a risky area that were hit by the flood (HitRisk) and homes that were not hit but are located in a risk zone of an affected municipality, and so can be considered a near-miss (NoHitRisk). [^27] We experimented alternative ways to characterize near-misses, including distance thresholds, with similar results. Our preferred assumption is the municipality because it does not require specifying a hard threshold for being considered a near-miss or not. This distinction is exemplified in Figure 2b.
@@ 3.3 Disaggregated impacts of flood risk
Our second research question aims at exploring the influence of flood risk ($R_{i}$) on home prices across all the national territory. Our baseline equation, Equation (2), is the starting point for this analysis. However, by interacting the indicator $R_{i}$ with another categorical/binary variable we can provide more insights about the heterogeneity beyond the global coefficient $\beta$.
First, we disaggregate the effects of buying a home at risk by buyers’ flood awareness at the time of the transaction, i.e. our measure of historical memory derived from historical flood records (Section 2.4). To this aim, we interact the risk indicator with a categorical indicator Awareness detailing the level of awareness in the region at the time of sale. As categories, we consider the terciles of the sale-time awareness distribution across all Italy, and name them high, medium and low awareness. This results in the following equation
$\log{P_{ikt}}=\alpha_{k}+\alpha_{t}+\rho R_{i}\times\texttt{Awareness}_{i}+\boldsymbol{\gamma}^{\prime}\textbf{X}_{i}+\epsilon_{ikt}.$ (4)
@@
Next, we also disaggregate the effects of buying a home at risk across specific regions, providing a more granular overview of the heterogeneous impacts of flood risk across Italy.
3.4 Disaggregation by socio-economic groups
To explore how buyers from different socio-economic groups may behave differently when buying homes at risk of flooding (research question 3), we build on the approach presented in Section 3.3 and focus on age and income as demographic variables.
We encoded the age of the buyer in the variable Age with two categories: Young and NotYoung. This is based on whether the mortgage was a mutuo giovani (Young Buyers’ Mortgage) or not. Transactions financed by the mutuo giovani are associated with buyers under 36 with an adjusted annual income lower than 40,000€. We assumed all mortgages not labeled in the data as mutuo giovani to refer to NotYoung buyers. [^28] Note that only individuals under the age of 36 are eligible to apply for a mutuo giovani. Conversely, while it is possible for a young buyer to obtain a standard mortgage, this is uncommon due to the more favorable terms offered by the mutuo giovani.
For the level of income of the buyers, we divided buyers into 3 levels of income (high, medium, and low) based on the three terciles of the income distribution for each region. We define income levels by region because average income is widely different across Northern and Southern regions, and we want this variable to capture relative income within regions.
We focus on two sub-research questions:
Question 3a: what discount do different groups of buyers obtain?
By performing the triple interaction between the $R_{i}$ indicator, the Awareness and Age variables, we can estimate whether young or more experienced buyers obtain different price reductions when buying homes at risk under different flood awareness levels. The equation used for this analysis is the following:
@@ 3.4 Disaggregation by socio-economic groups
$\log{P_{ikt}}=\alpha_{k}+\alpha_{t}+\zeta R_{i}\times\texttt{Awareness}_{i}\times\texttt{Age}_{i}+I+\boldsymbol{\gamma}^{\prime}\textbf{X}_{i}+\epsilon_{ikt}.$ (5)
In the above equation, the key coefficient is $\zeta$, capturing the effect of risk on home prices for each combination of age and awareness. In this regression we also included the term $I$ to control for buyers’ income, accounting for income differences between Young and NotYoung buyers (12%).
To further disaggregate this effect by the income levels of the buyer, we used the following variation of Equation (5):
@@
$\log{P_{ikt}}=\alpha_{k}+\alpha_{t}+\psi R_{i}\times\texttt{Awareness}_{i}\times\texttt{Age}_{i}\times\texttt{IncomeLevel}_{i}+I+\boldsymbol{\gamma}^{\prime}\textbf{X}_{i}+\epsilon_{ikt}$ (6)
@@
Question 3b: what is the income of the buyers moving to at-risk areas?
Apart from the discount they obtain, low-income or high-income buyers may be over-represented in at-risk areas. To explore this, we follow Zivin et al. (2023) and estimate the following regression:
$\log{I_{ikt}}=\alpha_{k}+\alpha_{t}+\beta R_{i}+\omega\texttt{Age}_{i}+\boldsymbol{\gamma}^{\prime}\textbf{X}_{i}+\epsilon_{ikt},$ (7)
where $\log{I_{ikt}}$ is the log of the household’s income, $\texttt{Age}_{i}$ is used as a dummy to control for young buyers (which have substantially lower income), and the other terms are equivalent to the ones described in Equation (2). In some specification, we also control for transaction price, in order to check if differences in income are mainly driven by cheaper homes or by alternative reasons. [^29] The price and income corrections implemented respectively in Equations (6) and (7), consider the variable with a logarithmic form.
@@ 4 Results
4.1 Specific floods only affect home prices in flooded areas
@@
To assess the impact that specific floods had on home prices we first focus on the major flood that hit Italy in the last 20 years, namely the May 2023 flood in Emilia-Romagna (coded EMSN154 by Copernicus). Because this flood is very recent, we only have observations up to 15 months after the flood event. We divide the year following the event in quarters and consider the interactions of quarters with HitRisk and NoHitRisk dummies (Equation 3). For instance, the interaction between a sale in the first quarter after the event and HitRisk indicates the reduction in price for homes that were hit by the flood and sold within 3 months of the event.
Figure 3 shows the results. The left panel shows the baseline discount for being in a risk zone for homes that were hit ($\beta_{1}$) and that were not hit ($\beta_{2}$). These coefficients state that HitRisk and NoHitRisk homes sell –on average– around 3% less than homes in safe areas, in line with results reported in the next subsection.
@@
These results were also consistent when considering alternative parameter specifications (such as more granular fixed effects), when analyzing only ground-floor apartments or independent houses, and when exploring other major flood events (see Supplementary Section B.1). [^33] Our study on the 2023 Emilia-Romagna flood (EMS154) can be consistently compared to previous studies focusing on homes in the 100-year floodplain (Atreya et al., 2013; Bin and Landry, 2013). In fact, almost all at-risk homes in the municipalities affected by the Emilia-Romagna flood are in high or medium risk areas; thus, they are characterized by return times less or equal than 200 years. We only find a negligible number of homes placed in the low-risk zones (9 homes in total). To the best of our knowledge, considering ISPRA data, high or medium risk areas are the those that better align with the 100-years flood zones used in previous studies.
@@ 4.2 Homes at risk sell at lower price under high awareness conditions
In the left panel of Figure 3, we showed that being located in an at-risk area leads to a lower home price, irrespective of the specific flood considered. To explore this question further, we estimate Equation (2) on our entire dataset aiming to provide a national-scale picture of flood risk across Italy.
@@
Model 4 shows the disaggregation of flood risk effects when considering our awareness measure at the time of sale. Notably, the statistically significant 1% price reduction at the national scale is mainly driven by the discount observed in high awareness conditions (-1.7%). The other coefficients for flood risk are not statistically significant and show an average 0.6% and 0.2% price reduction for transactions with medium and low awareness. The leading role of high awareness transactions in determining the impacts of flood risk on prices is robust considering more granular fixed effects (Model 6 of Table 2), different event data and half-life parameter (we refer the reader to Supplementary Section B.2, Tables A7 and A8, for further details).
@@
4.3 Differential effects across socio-economic groups
Experienced and higher-income buyers obtain stronger price reductions
The results presented in the previous sections support our hypothesis that flood-risk-driven home price reductions are due to the historical memory of previous floods and not to specific events. Here, we investigate whether buyers with different socio-economic status (i.e., age and income) obtain different discounts when buying at-risk homes.
We start by analyzing the differences in home prices considering buyers’ age (as proxied by under-36 mutuo giovani, see Section 3.4). Figure 5 shows the discount obtained by Young and NotYoung buyers when purchasing homes at risk, distinguishing by the level of flood awareness (high, medium and low) at the time of sale. Under low awareness conditions, Young and NotYoung buyers do not get any statistically significant discount for homes in flood-prone areas. As flood awareness gets higher, NotYoung buyers exhibit higher and higher discounts, reaching 2% for transactions of risky homes in high awareness times. Conversely, we do not find a clear trend for Young buyers, with no statistically significant discounts under any awareness level. We interpret this result as evidence that experienced buyers, who may have witnessed several flood events in their lifetime, pay more attention to flood risk when purchasing a home. This is because other differences between buyers in their ability to negotiate discounts are unlikely to vary so markedly based on flood risk awareness.
Figure 5: Flood risk effect on home prices disaggregated across flood awareness (rows) and mortgage applicant’s age, Equation (5). The age is proxied considering whether the transaction is financed by a mutuo giovani (red dot) or not (blue square).
Building on the results presented in Figure 5, we deepen our analysis by considering also the income class of the buyers, following Equation (6). Figure 6 shows our findings. Though on average Young buyers exhibit no significant price reduction for buying homes at risk, we observe that this behavior changes when disaggregating by income levels. Young buyers (left panel) with high-income obtain a statistically significant price reduction (around 3.5%) when purchasing at-risk homes under high and low awareness conditions. This behavior is not observed in the other income levels, as low- and medium-income buyers purchase homes in flood-prone areas at a premium, ranging from 0.1-2% depending on the awareness level. The influence of income is not so marked for the NotYoung buyers (right panel). In this case, all income classes have a similar trend that follows the income-aggregated average from Figure 5. Interestingly, all income levels achieve a statistically significant discount under high awareness conditions (around 2.5%), but this effect vanishes with low awareness.
Overall, NotYoung buyers seem to be more influenced by the level of awareness all income levels than Young buyers. This further supports our hypothesis that flood risk is engraved in the historical exposure to prior floods. Conversely, other socio-economical factors, like income, are key determinants for transactions performed by Young buyers. In fact, only Young buyers with high-income purchase homes at risk with significant price reductions, while other income classes pay a premium for risky homes. Further research is required to explore the nexus among income, education, and perception of flood risk (and more in general climate-related risks). However, we leave the exploration of this hypothesis to future work since it requires additional data (such as the precise education level of the buyers) that we lack.
@@ 4.3 Differential effects across socio-economic groups
Low-income buyers are more likely to move to at-risk areas
Having shown that buyers’ income plays a key role in determining the effect of flood risk on home prices, we now want to check if households that buy homes in at-risk areas have lower or higher incomes than households purchasing homes elsewhere. Households that buy in at-risk areas have a mean monthly income of 2,898€, while households that buy elsewhere have a mean monthly income of 2,953€ (1.9% larger). However, this unconditional comparison is likely to be biased if at-risk areas are concentrated in the richest parts of Italy, such as the Po Valley.
Table 3 reports the effect of flood risk on buyers’ income, estimated following Equation (7). We study the risk coefficient on incomes, even including the transaction price as an independent variable. If including the price as a dependent variable makes any effect on income non-significant, it suggests that lower prices are sufficient to explain why lower-income households sort into at-risk areas. Vice versa, other factors such as a lack of information on flood risk for lower-income households may be at play.
Table 3: Flood risk effect on buyers’ incomes, Equation (7). Models 1-2 report the effect of the risk on buyers’ income at the national scale (Model 1) and disaggregated by the level of flood awareness at the time of the purchase (Model 2, with half life time equal to 10 years). Models 3-4 show the same analyses, including the logarithm of the home sale price as an independent variable. *** $p<0.01$, ** $p<0.05$, * $p<0.1$.

### v3

@@ 3 Econometric design
To quantify the influence of flood risk on the Italian housing market, we have addressed three separate research questions: (1) Do we observe changes in home prices after specific flood events? (2) Does flood risk affect home prices in general? Is the effect larger by repeated exposure to floods (awareness) or by socioeconomic characteristics of the buyers? (3) Are the incomes of buyers purchasing properties in flood risk areas different from the income of buyers in areas that are not at risk? In the following, we describe how we address all these questions by building on a baseline hedonic regression.
@@ 3.1 Baseline hedonic regression
In our formulation, $\alpha_{k}$ captures all time-invariant home price differences due to being located in spatial unit $k$. We consider different definitions of spatial units, at different spatial aggregation levels. The widest definition of $k$ is the municipality, with a median area of 22 km2. Especially for large municipalities, it is unlikely that this fixed effect captures enough difference in location, so we consider two further narrower definitions of $\alpha_{k}$. These are the OMI microzone and the census tract, respectively featuring a median area of 1.5 km2 and 0.03km2.242424The OMI microzone is the standard spatial unit of the Osservatorio del Mercato Immobiliare (OMI), the housing market observatory of the Italian Tax Office. It roughly corresponds to a neighborhood, although in rural areas it may contain larger portions of territory. Instead, census tracts correspond to blocks in cities and to villages (“frazioni”) in rural areas. We use the latest available census tracts, namely from the 2011 census. The inclusion of the $\alpha_{k}$ term completely captures the spatial correlation existing in our data, see Supplementary Section B.2. This confirms that our hedonic regression approach is adequate and spatial econometric approaches are not necessary.
@@
3.2 Effect of single flood events
To estimate the impact that specific flood events had on home price (research question 1) we built on Equation (2) to formulate a hedonic difference-in-difference specification. Following Ortega and Taṣpınar (2018), we disaggregated the $R_{i}$ indicator into two categories, homes in a risky area that were hit by the flood (HitRisk) and homes that were not hit but are located in a risk zone of an affected municipality, and so can be considered a near-miss (NoHitRisk). [^26] We experimented alternative ways to characterize near-misses, including distance thresholds, with similar results. Our preferred assumption is the municipality because it does not require specifying a hard threshold for being considered a near-miss or not. This distinction is exemplified in Figure 2b. [^27] Homes hit by the Emilia-Romagna flood (thus, sold after the flood), but not at risk (HitNoRisk) were not considered in the analyses because we have only 3 homes in this situation (we remind that the homes in the affected municipalities are 14977 and those hit by the flood are 1464, see Table A2).
@@ 3.3 Disaggregated impacts of flood risk
Our second research question aims at exploring the influence of flood risk ($R_{i}$) on home prices across all the national territory depending on different awareness conditions and buyers’ characteristics. Our baseline equation, Equation (2), is the starting point for this analysis. However, by interacting the indicator $R_{i}$ with other categorical/binary variables we can provide more insights about the heterogeneity beyond the global coefficient $\beta$.
Disaggregation by region.
@@
We disaggregated the effects of buying a home at risk across specific regions, providing a more granular overview of the heterogeneous impacts of flood risk across Italy. To do so we applied the following regression
@@
$\log{P_{ikt}}=\alpha_{k}+\alpha_{t}+\omega R_{i}\times\texttt{Region}_{i}+\boldsymbol{\gamma}^{\prime}\textbf{X}_{i}+\epsilon_{ikt}.$ (4)
Disaggregation by awareness.
We also disaggregated the effects of buying a home at risk by buyers’ flood awareness at the time of the transaction, i.e. our measure of historical memory derived from historical flood records (Section 2.4). To this aim, we interact the risk indicator with a categorical indicator Awareness detailing the level of awareness in the region at the time of sale. As categories, we consider the terciles of the sale-time awareness distribution across all Italy, and name them high, medium and low awareness [^28] Categorical terms were introduced as dummy variables. Thus, when not interacted, we choose to remove the low awareness dummy to prevent colinearity.. This results in the following equation
@@
$\log{P_{ikt}}=\alpha_{k}+\alpha_{t}+\rho R_{i}\times\texttt{Awareness}_{i}+\beta_{aw}\texttt{Awareness}_{i}+\boldsymbol{\gamma}^{\prime}\textbf{X}_{i}+\epsilon_{ikt}.$ (5)
Disaggregation by socio-economic groups.
Although the theoretical foundations of hedonic regressions (Rosen, 1974) rely on competitive markets in which the marginal buyer alone determines prices, real housing markets are far from this ideal. Strong segmentation (Piazzesi et al., 2020), together with liquidity constraints and information asymmetries, allows for systematic price discrimination across different types of buyers. Following Bernstein et al. (2019), we investigate whether buyers differ in their willingness to pay for flood risk depending on their age and income.
We encode the age of the buyer in the variable Age with two categories: Young and NotYoung. This is based on whether the mortgage was a mutuo giovani (Young Buyers’ Mortgage) or not. Transactions financed by the mutuo giovani are associated with buyers under 36 with an adjusted annual income lower than 40,000€. We assumed all mortgages not labeled in the data as mutuo giovani to refer to NotYoung buyers. [^29] Note that only individuals under the age of 36 are eligible to apply for a mutuo giovani. Conversely, while it is possible for a young buyer to obtain a standard mortgage, this is uncommon due to the more favorable terms offered by the mutuo giovani.
For the level of income of the buyers, we divided buyers into 3 levels of income (high, medium, and low) based on the three terciles of the income distribution for each region. We define the IncomeLevel of the buyers of each transaction by region because average income is widely different across Northern and Southern regions, and we want this variable to capture relative income within regions.
By interacting the $R_{i}$ indicator, Awareness, Age, and IncomeLevel, we can estimate whether young or more experienced buyers in different income classes obtain different price reductions when buying homes at risk across flood awareness levels. The equation used for this analysis is the following:
$\log{P_{ikt}}=$ $\alpha_{k}+\alpha_{t}+\psi R_{i}\times\texttt{Awareness}_{i}\times\texttt{Age}_{i}\times\texttt{IncomeLevel}_{i}+$
$\quad\beta_{in}\texttt{IncomeLevel}_{i}+\beta_{age}\texttt{Age}_{i}+\beta_{aw}\texttt{Awareness}_{i}+\boldsymbol{\gamma}^{\prime}\textbf{X}_{i}+\epsilon_{ikt},$ (6)
@@
3.4 Income of the buyers moving to at-risk areas
Beyond the evaluation of home prices in flood prone areas, we now focus on the income of people settling in risky areas. To explore this, we follow Zivin et al. (2023) and estimate the following regression:
$\log{I_{ikt}}=\alpha_{k}+\alpha_{t}+\beta R_{i}+\boldsymbol{\gamma}^{\prime}\textbf{X}_{i}+\epsilon_{ikt},$ (7)
where $\log{I_{ikt}}$ is the log of the household’s income and the other terms are equivalent to the ones described in Equation (2).
Next, following the analyses on home prices, we zoom into the coefficient $\beta$ in Eq. (7) to explore whether flood-risk awareness and buyers’ age play a role. For awareness, we change Eq. (5) to consider income ($I$) as dependent variable. For age, we consider the following specification:
$\log{I_{ikt}}=$ $\alpha_{k}+\alpha_{t}+\phi R_{i}\times\texttt{Age}_{i}\times\texttt{Awareness}_{i}+$
$\quad\beta_{age}\texttt{Age}_{i}+\beta_{aw}\texttt{Awareness}_{i}+\boldsymbol{\gamma}^{\prime}\textbf{X}_{i}+\epsilon_{ikt},$ (8)
where the coefficient $\phi$ quantifies whether young or experienced buyers purchasing homes at risk in certain awareness conditions have higher or lower income compared to similar people in areas that are not at risk.
@@ 4 Results
4.1 Single floods only affect home prices in flooded areas
@@
To assess the impact that specific floods had on home prices we first focus on the major flood that hit Italy in the last 20 years, namely the May 2023 flood in Emilia-Romagna (coded EMSN154 by Copernicus). We later provide similar analyses for the 2016 Piemonte-Liguria (code EMSR192) and 2023 Toscana (EMSR705) floods in Supplementary Section B.1.
Since the Emilia-Romagna flood is very recent, we only have observations up to 15 months after the flood event. We divide the year following the event in quarters and consider the interactions of quarters with HitRisk and NoHitRisk dummies (Equation 3). For instance, the interaction between a sale in the first quarter after the event and HitRisk indicates the reduction in price for homes that were hit by the flood and sold within 3 months of the event.
Figure 3 shows the results. The left panel shows the baseline discount for being in a risk zone for homes that were hit ($\beta_{1}$) and that were not hit ($\beta_{2}$). These coefficients indicate that HitRisk and NoHitRisk homes sell –on average– around 3% less than homes in safe areas, in line with results reported in the next subsection.
@@
Table A3 compares the statistical properties of the homes considered one year before and one year after this flood. These results show that the hedonic price function did not change after the flood event, supporting our interpretation of the coefficients. These results were also consistent when considering alternative parameter specifications (such as more granular fixed effects), when analyzing only ground-floor apartments or independent houses, and when exploring other major flood events (see Supplementary Section B.1). [^33] Our study on the 2023 Emilia-Romagna flood (EMS154) can be consistently compared to previous studies focusing on homes in the 100-year floodplain (Atreya et al., 2013; Bin and Landry, 2013). In fact, almost all at-risk homes in the municipalities affected by the Emilia-Romagna flood are in high or medium risk areas; thus, they are characterized by return times less or equal than 200 years. We only find a negligible number of homes placed in the low-risk zones (9 homes in total). To the best of our knowledge, considering ISPRA data, high or medium risk areas are the those that better align with the 100-years flood zones used in previous studies. See also footnote 18.
@@ 4.2 Homes at risk sell at lower price under high awareness conditions
In the left panel of Figure 3, we showed that being located in an area at risk leads to a lower home price, irrespective of the specific flood considered. To explore this question further, we estimate Equation (2) on our entire dataset aiming to provide a national-scale picture of flood risk across Italy.
@@
Model 4 shows the disaggregation of flood risk effects when considering our awareness measure at the time of sale. Notably, the statistically significant 1% price reduction at the national scale is mainly driven by the discount observed in high awareness conditions (-1.7%). The other coefficients for flood risk are slightly or not statistically significant and show an average 0.7% and 0.2% price reduction for transactions with medium and low awareness. The leading role of high awareness transactions in determining the impacts of flood risk on prices is robust considering more granular fixed effects (Model 6 of Table 2), different event data and half-life parameter (we refer the reader to Supplementary Section B.2, Tables A7 and A8, for further details).
@@
Higher-income buyers obtain stronger price reductions
The results presented in the previous sections support our hypothesis that flood-risk-related price discounts are driven by the historical memory of past floods rather than by the occurrence of specific flood events. In this section, we investigate whether buyers with different socio-economic characteristics (specifically age and income) obtain heterogeneous discounts when purchasing homes exposed to flood risk.
To this end, we estimate Eq. (6) (see Section 3.3). Figure 5 reports the estimated flood-risk discount for Young and NotYoung buyers purchasing at-risk properties, distinguishing by buyers’ income levels and by the degree of flood-risk awareness (high, medium, and low). Under high-awareness conditions, Young (left panel) and NotYoung (right panel) buyers display remarkably similar patterns: high-income buyers obtain a flood-risk discount of approximately 3%, while medium-income buyers receive a discount of around 2%. As flood-risk awareness declines, the effect of flood exposure becomes statistically insignificant across most groups. The only notable exception is Young high-income buyers in low-awareness areas, who still obtain a discount of roughly 2% when purchasing homes at risk.
Taken together, these findings indicate that income, rather than age, is the main driver of heterogeneity in the price response to flood risk. While a full interpretation is beyond the scope of this paper, the observed patterns may reflect differences in financial constraints, information acquisition, education, or bargaining power across income groups.
@@
4.3 Low-income buyers are more likely to move to at-risk areas
We now examine whether households purchasing homes in flood-risk areas differ systematically in income from those purchasing properties elsewhere, addressing our final research question. A simple comparison of group means shows that households buying in at-risk areas have an average monthly income of 2,898€, compared to 2,953€for households buying in non-risk areas—a difference of about 1.9%. However, this unconditional comparison may be misleading, as flood-risk exposure is not randomly distributed across space and is more prevalent in relatively affluent regions of Italy, such as the Po Valley.
Table 3 reports estimates of the effect of flood risk on buyers’ income using Equation (7). Focusing on the coefficient on flood risk, Model 1 shows that, on average across Italy, buyers purchasing properties in flood-risk areas have incomes approximately 0.9% lower than those purchasing in non-risk areas. Model 2 reveals substantial heterogeneity across awareness levels: income differences are larger and statistically significant in high- and medium-awareness contexts (–1.4% and –1.2%, respectively), while the effect becomes small and statistically insignificant under low-awareness conditions.
We further explore heterogeneity by buyers’ age by estimating Equation (8). The results, shown in Figure 6, indicate pronounced age-related differences. Among NotYoung buyers, incomes of those purchasing homes in flood-risk areas are approximately 3% lower than the incomes of buyers that do not purchase properties at risk in high-awareness settings. While this income gap narrows as awareness declines, it remains negative across all awareness levels. In contrast, results for Young buyers are less consistent and in several cases not statistically significant.
Overall, these findings suggest that income-based sorting into flood-risk areas is particularly pronounced among older households. This pattern is likely shaped by differences in financial constraints, information, and bargaining power across demographic groups, consistent with the heterogeneity found in Section 4.2.
Table 3: Flood risk effect on buyers’ incomes, Equation (7). Models 1-2 report the effect of the risk on buyers’ income at the national scale (Model 1) and disaggregated by the level of flood awareness at the time of the purchase (Model 2, with half life time equal to 10 years). *** $p<0.01$, ** $p<0.05$, * $p<0.1$.